ganeshie8 (ganeshie8):

If \(a\) and \(b\) are the roots of \(x^2-2x+4=0\), then \(a^9 +b^9 = ?\\ a^{10} + b^{10} = ?\)

OpenStudy (perl):

you can use demoivre's theorem , if you change the complex roots to polar trig form

OpenStudy (perl):

or exponential trig form

ganeshie8 (ganeshie8):

that works nicely!!!

OpenStudy (studygurl14):

It makes sense because of the \(\large \pm\) in the quadratic formula.

OpenStudy (studygurl14):

To me anyway. That logic could be completely wrong...but, idk

OpenStudy (dan815):

a,b= x^2-2x-1+4=0 (x-1)^2=-3 x=+/- sqrt3 i + 1 \[a= 1+\sqrt3 i\\ b= 1-\sqrt3 i\] polar form now

OpenStudy (dan815):